OpenStudy (anonymous):

Solve the equation (eqn coming soon) . . ANSWER = {-1,5 pm 2 sqrt6

OpenStudy (anonymous):

\[(x-\frac{ 1 }{ x })^2-8(x+\frac{ 1 }{x })-16=0\]

OpenStudy (anonymous):

i have solved this and i am getting 2 values of y

OpenStudy (anonymous):

\[y=4+4\sqrt{2}, y=4-4\sqrt{2}\]

hartnn (hartnn):

have u written (x-1/x)^2 as (x+1/x)^2 -4 ? and then put y= x+1/x

OpenStudy (anonymous):

yes i have done it all

hartnn (hartnn):

then u should get y =10 or -2

OpenStudy (anonymous):

i asked a question like this hours ago right

hartnn (hartnn):

yes. and i have helped u solve..... why u not getting y = 10,-2 ??

OpenStudy (anonymous):

y^2-8y-16=0 this is the quadratic

hartnn (hartnn):

\(\huge (x-1/x)^2 as (x+1/x)^2 -4\)

hartnn (hartnn):

u forgot -4

OpenStudy (anonymous):

y^2-8y-20=0 is'nt it

hartnn (hartnn):

yes. y=?

OpenStudy (anonymous):

y=10 and y=-2

hartnn (hartnn):

yesm now put y=x+1/x

OpenStudy (anonymous):

10=x+1/x right?

hartnn (hartnn):

yes, solve that quadratic.....

OpenStudy (anonymous):

done

OpenStudy (anonymous):

\[x=5\pm 4\sqrt{6} and x=-1\]